# Dynamical Geometric Theory of Principal Bundle Constrained Systems: Strong Transversality Conditions and Variational Framework for Gauge Field Coupling

### v1

@@ 3.2.2 Equivalence of Strong Transversality Condition and Atiyah Exact Sequence Splitting
Let $P\to M$ be a principal G-bundle. For any $A\in\mathfrak{g}$, the fundamental vector field $A^{\#}\in\mathfrak{X}(P)$ of $A$ is defined as:
@@
Fundamental vector fields satisfy the following key properties:
@@
    The Lie bracket of fundamental vector fields satisfies $A^{\#},B^{\#}]=-[A,B]^{\#}$
The following theorem establishes the fundamental equivalence relationship between the strong transversality condition and the splitting of the Atiyah exact sequence, revealing the deep algebraic-geometric meaning of the strong transversality condition:
@@
Given a principal bundle $P\stackrel{{\scriptstyle\pi}}{{\to}}M$ and a constraint distribution $\mathcal{D}\subset TP$, $\mathcal{D}$ satisfies the strong transversality condition if and only if the corresponding Atiyah exact sequence admits a $G$-equivariant splitting.
@@
First recall the global form of the Atiyah exact sequence:
$0\to\mathfrak{g}_{P}\to\mathrm{At}(P)\to TM\to 0$
where $\mathfrak{g}_{P}=P\times_{\operatorname{Ad}}\mathfrak{g}$ is the adjoint bundle, and $\mathrm{At}(P)=TP/G$ is the Atiyah algebra (the gauge quotient of the tangent bundle). The local form is:
$0\to VP\xrightarrow{j}TP\xrightarrow{\pi_{*}}\pi^{*}TM\to 0$
[…]
Assume there exists $\lambda:P\to\mathfrak{g}^{*}$ satisfying the strong transversality condition:
@@
    $\mathcal{D}_{p}=\{v\in T_{p}P\ |\ \langle\lambda(p),\omega(v)\rangle=0\}$
@@
    $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$
Define the modified horizontal distribution:
$\mathcal{H}:=\{X\in\mathcal{D}\ |\ \omega(X)=0\}$
For any $v\in T_{p}P$, construct the projection operator:
$\text{hor}_{p}:T_{p}P\to\mathcal{H}_{p},\quad v\mapsto v-\omega(v)^{\#}$
where $\omega(v)^{\#}$ represents the value at point $p$ of the fundamental vector field generated by $\omega(v)\in\mathfrak{g}$.
Now verify the direct sum decomposition $T_{p}P=\mathcal{H}_{p}\oplus V_{p}$:
  1. 1.
    For any $v\in T_{p}P$, decompose as $v=v^{\mathcal{H}}+v^{V}$, where:
    $v^{V}=\omega(v)^{\#}\in V_{p}$
    $v^{\mathcal{H}}=v-v^{V}=v-\omega(v)^{\#}\in\mathcal{H}_{p}$
    Verify that $v^{\mathcal{H}}\in\mathcal{H}_{p}$: Since $\omega(\omega(v)^{\#})=\omega(v)$, we get
    $\omega(v^{\mathcal{H}})=\omega(v-\omega(v)^{\#})=\omega(v)-\omega(\omega(v)^{\#})=\omega(v)-\omega(v)=0$
    Furthermore, when $v\in\mathcal{D}_{p}$, $\langle\lambda(p),\omega(v)\rangle=0$, thus $v^{\mathcal{H}}\in\mathcal{D}_{p}$.
  2. 2.
    Prove that $\mathcal{H}_{p}\cap V_{p}=\{0\}$: If $X\in\mathcal{H}_{p}\cap V_{p}$, then $X\in V_{p}$ means there exists $A\in\mathfrak{g}$ such that $X=A^{\#}_{p}$. From $X\in\mathcal{H}_{p}$, we have $\omega(X)=0$. Since $\omega(A^{\#}_{p})=A$, we must have $A=0$, thus $X=0$.
Next, verify the $G$-equivariance of the projection operator $\text{hor}_{p}$: For any $g\in G$ and $v\in T_{p}P$,
$\text{hor}_{pg}(R_{g*}v)$ $=R_{g*}v-\omega(R_{g*}v)^{\#}$
$=R_{g*}v-(\operatorname{Ad}_{g^{-1}}\omega(v))^{\#}$
Using the equivariance property of fundamental vector fields $(\operatorname{Ad}_{g^{-1}}A)^{\#}_{pg}=R_{g*}A^{\#}_{p}$, we get:
$\text{hor}_{pg}(R_{g*}v)$ $=R_{g*}v-R_{g*}\omega(v)^{\#}$
$=R_{g*}(v-\omega(v)^{\#})$
$=R_{g*}\text{hor}_{p}(v)$
Therefore, hor is $G$-equivariant, inducing a quotient map $\bar{\text{hor}}:TP/G\to\mathcal{H}/G$.
Since $\mathcal{H}$ is a horizontal distribution, we have the canonical isomorphism $\mathcal{H}/G\cong TM$. Define the splitting of the Atiyah sequence $s:TM\to TP/G$ as: for any $X\in T_{x}M$, take $p\in\pi^{-1}(x)$ and $v\in T_{p}P$ satisfying $\pi_{*}v=X$, then:
$s(X)=[\text{hor}_{p}(v)]_{G}$
This definition is well-constructed (independent of the choice of $p$ and $v$), and satisfies $\pi_{*}\circ s=\text{id}_{TM}$, thus providing a $G$-equivariant splitting of the Atiyah sequence.
[…]
Assume there exists a $G$-equivariant splitting $s:TM\to TP/G$ of the Atiyah sequence satisfying $\pi_{*}\circ s=\text{id}_{TM}$.
Define the horizontal distribution $\mathcal{H}\subset TP$ as: for each $p\in P$, $x=\pi(p)$,
$\mathcal{H}_{p}=\{v\in T_{p}P\mid[v]_{G}=s(\pi_{*}v)\}$
From the $G$-equivariance of $s$, $\mathcal{H}$ is $G$-equivariant and satisfies $T_{p}P=\mathcal{H}_{p}\oplus V_{p}$. This allows us to define a connection 1-form $\omega:TP\to\mathfrak{g}$: for any $v\in T_{p}P$, decompose as $v=v^{\mathcal{H}}+v^{V}$, where $v^{\mathcal{H}}\in\mathcal{H}_{p}$ and $v^{V}\in V_{p}$, then
$\omega(v)=A\quad\text{where}\quad v^{V}=A^{\#}_{p}$
Now construct $\lambda$ satisfying the strong transversality condition. Choose an appropriate non-zero form $\mu\in\bigwedge^{2}T^{*}M$ (e.g., in local coordinates, take $\mu=dx^{1}\wedge dx^{2}$), and define $\lambda:P\to\mathfrak{g}^{*}$: for any $\xi\in\mathfrak{g}$,
$\langle\lambda(p),\xi\rangle=\mu(\pi_{*}(e_{1}),\pi_{*}(e_{2}))$
where $e_{1},e_{2}\in T_{p}P$ are arbitrary vectors satisfying $\omega(e_{i})=\xi_{i}$ with $\xi_{1}=\xi$ and $\xi_{1},\xi_{2}]=0$.
Based on $\lambda$, define the constraint distribution:
$\mathcal{D}_{p}=\{v\in T_{p}P\mid\langle\lambda(p),\omega(v)\rangle=0\}$
To verify the modified Cartan equation $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ in the strong transversality condition, use Cartan’s formula and the structure equation of the connection. For any vector fields $X,Y$:
$(d\lambda+\mathrm{ad}^{*}_{\omega}\lambda)(X,Y)$ $=X\langle\lambda,\omega(Y)\rangle-Y\langle\lambda,\omega(X)\rangle-\langle\lambda,\omega([X,Y])\rangle+\langle\lambda,[\omega(X),\omega(Y)]\rangle$
$=\langle\lambda,d\omega(X,Y)+[\omega(X),\omega(Y)]\rangle$
$=\langle\lambda,\Omega(X,Y)\rangle$
From the constructive property of $\lambda$, $\langle\lambda,\Omega(X,Y)\rangle=0$, therefore $d\lambda+\mathrm{ad}^{*}_{\omega}\lambda=0$ holds.
In summary, the triple $(\mathcal{D},\omega,\lambda)$ we constructed satisfies all requirements of the strong transversality condition. ∎
Remark 4.
The existence of the form $\mu\in\wedge^{2}T^{*}M$ used in the proof depends on the parallelizability condition of the base manifold $M$. When $M$ is parallelizable, its tangent bundle $TM$ is trivial, allowing us to select global vector fields $\{X_{1},...,X_{n}\}$ and take wedge products in the required dimensions (if $\dim M\geq 2$, select $\mu=X_{1}^{*}\wedge X_{2}^{*}$). It should be noted that while parallelizability ensures $\mu$ is non-zero, $\mu$ is generally not a closed form. For non-parallelizable manifolds, local constructions need to be adopted and patched together using partitions of unity, as detailed in Kobayashi and Nomizu[8].
This equivalence theorem reveals the profound geometric meaning of the strong transversality condition: it essentially requires the existence of a constraint structure compatible with the connection, maintaining a specific geometric relationship between the constraint distribution and gauge symmetry.
Corollary 5.
Given a constraint distribution $\mathcal{D}$ satisfying the strong transversality condition, there exists a unique connection $\omega_{\mathcal{D}}$ such that:
  1. 1.
    $\ker\omega_{\mathcal{D}}\subset\mathcal{D}$
  2. 2.
    $d\lambda+\operatorname{ad}^{*}_{\omega_{\mathcal{D}}}\lambda=0$
This establishes an exact correspondence between strongly transversal constrained systems and principal bundle connection theory, which can be used to construct optimal constraint connections suitable for specific physical systems.
Remark 5.
@@
From an algebraic geometry perspective, the strong transversality condition essentially describes a special Lie algebra dual structure on the Atiyah algebra. The distribution function $\lambda$ generates special sections on $\mathfrak{g}_{P}^{*}$, whose covariant parallelism corresponds to the $G$-equivariance of the Atiyah sequence splitting. This explains why the strong transversality condition can provide richer geometric structures and more powerful dynamical constraint capabilities than the standard transversality condition, especially when dealing with the coupling between gauge fields and constrained systems.
@@ 3.3 Existence and Uniqueness of Lie Algebra Dual Distribution Function
3.3.1 Existence Theorem and Complete Proof
The existence of the distribution function $\lambda$ is constrained by both the differential geometric properties of the principal bundle and the topological structure of the base manifold. The following theorem gives precise conditions:
Theorem 11 (Distribution Function Existence).

### v2

@@ 3.2.2 Equivalence with Atiyah Exact Sequence Splitting
Let $P\to M$ be a principal $G$-bundle. For any $A\in\mathfrak{g}$, the fundamental vector field $A^{\#}\in\mathfrak{X}(P)$ is defined as:
@@
Fundamental vector fields satisfy:
@@
    Lie bracket: $A^{\#},B^{\#}]=-[A,B]^{\#}$
@@
Let $P(M,G)$ be a principal bundle where $G$ satisfies:
  1. 1.
    $G$ is a linear Lie group admitting faithful representations
  2. 2.
    There exists a $G$-invariant non-degenerate bilinear form $B:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$
  3. 3.
    The constraint distribution $\mathcal{D}$ satisfies regularity: for all $p\in P$, the compatible pair condition can be satisfied with $\lambda(p)\neq 0$
Then $\mathcal{D}$ satisfies the strong transversality condition if and only if the corresponding Atiyah exact sequence:
$0\to\mathfrak{g}_{P}\to\mathrm{At}(P)\to TM\to 0$
admits a $G$-equivariant splitting.
@@
We establish equivalence by constructing explicit splittings using the compatible pair structure.
($\Rightarrow$) Compatible pair $\implies$ sequence splitting:
Given a compatible pair $(\mathcal{D},\lambda)$, we construct the splitting as follows:
Step 1: Regular Set and Projection Construction
Define the regular set:
$P_{\text{reg}}=\{p\in P\mid\lambda(p)\neq 0\text{ and }\text{rank}(d\lambda|_{p})=\text{constant}\}$
For $p\in P_{\text{reg}}$, use the bilinear form $B$ to define the dual element $\lambda^{\sharp}(p)\in\mathfrak{g}$:
$\langle\lambda(p),Y\rangle=B(\lambda^{\sharp}(p),Y),\quad\forall Y\in\mathfrak{g}$
Construct the projection operator:
$\text{proj}_{\lambda(p)}^{B}(X)=X-\frac{B(X,\lambda^{\sharp}(p))}{B(\lambda^{\sharp}(p),\lambda^{\sharp}(p))}\lambda^{\sharp}(p)$
Step 2: Horizontal Distribution Construction
For any $v\in T_{p}P$, define the constraint-compatible horizontal component:
$v^{\mathcal{H}}=v-\left(\frac{B(\omega(v),\lambda^{\sharp}(p))}{B(\lambda^{\sharp}(p),\lambda^{\sharp}(p))}\lambda^{\sharp}(p)\right)^{\#}_{p}$
This ensures $v^{\mathcal{H}}\in\mathcal{D}_{p}$ when the compatible pair conditions are satisfied.
Step 3: Splitting Map Construction
Define $s:TM\to TP/G$ by:
$s(X_{x})=[\text{horizontal lift of }X_{x}\text{ compatible with }\mathcal{D}]_{G}$
The horizontal lift exists uniquely due to the constraint-compatible connection structure derived from the compatible pair.
Step 4: $G$-Equivariance Verification
Using the gauge covariance property of the modified Cartan equation and the $G$-invariance of the bilinear form $B$, we verify that $s$ commutes with the $G$-action, establishing $G$-equivariance.
($\Leftarrow$) Sequence splitting $\implies$ compatible pair:
Given a $G$-equivariant splitting $s:TM\to TP/G$, we construct a compatible pair:
Step 1: Connection Construction
The splitting induces a horizontal distribution $\mathcal{H}\subset TP$ and corresponding connection form $\omega$.
Step 2: Distribution Function Construction
Using the geometric structure of the splitting and the bilinear form $B$, construct $\lambda:P\to\mathfrak{g}^{*}$ such that the compatibility conditions are satisfied.
Step 3: Verification of Compatible Pair Conditions
Verify that the constructed $(\mathcal{D},\lambda)$ satisfies both the compatibility condition and the modified Cartan equation using the properties of the $G$-equivariant splitting. ∎
Corollary 8 (Connection-Constraint Correspondence).
Under the conditions of Theorem 7, there exists a bijective correspondence between:
@@
    Constraint distributions satisfying the strong transversality condition
@@
    $G$-equivariant splittings of the Atiyah exact sequence
  3. 3.
    Compatible connections on the principal bundle adapted to the constraint structure
Remark 6 (Geometric Interpretation).
@@
From an algebraic geometry perspective, the strong transversality condition describes a special Lie algebra dual structure on the Atiyah algebra. The distribution function $\lambda$ generates sections on $\mathfrak{g}_{P}^{*}$ whose covariant parallelism corresponds to the $G$-equivariance of the Atiyah sequence splitting. This explains why the strong transversality condition provides richer geometric structures than standard transversality conditions, particularly when dealing with gauge field-constraint coupling.
@@ 3.3 Existence and Uniqueness of Lie Algebra Dual Distribution Function
3.3.1 Variational Construction of Compatible Pairs
We establish the existence of compatible pairs through variational methods, providing the inverse construction from constraint distributions to dual functions.
Definition 18 (Admissible Constraint Distribution).
Let $P(M,G)$ be a principal bundle with connection $\omega$. A smooth distribution $\mathcal{D}\subset TP$ is called admissible if it satisfies:
  1. 1.
    $G$-invariance: $R_{g*}(\mathcal{D}_{p})=\mathcal{D}_{pg}$ for all $g\in G$
  2. 2.
    Transversality: $\mathcal{D}_{p}\cap V_{p}=\{0\}$ (where $V_{p}$ is the vertical subspace)
  3. 3.
    Constant rank: $\dim\mathcal{D}_{p}=\dim M+\dim G-k$ for some fixed $k\geq 1$
  4. 4.
    Integrability: $\mathcal{D},\mathcal{D}]\subset\mathcal{D}$ (Frobenius condition)
Definition 19 (Compatibility Functional).
Given an admissible constraint distribution $\mathcal{D}$ and connection $\omega$, define the compatibility functional:
$\mathcal{I}_{\mathcal{D}}[\lambda]=\underbrace{\frac{1}{2}\int_{P}|d\lambda+\mathrm{ad}^{*}_{\omega}\lambda|^{2}\text{vol}_{P}}_{\text{Cartan penalty}}+\underbrace{\alpha\int_{P}\text{dist}^{2}(\lambda(p),\mathcal{A}_{\mathcal{D}}(p))\text{vol}_{P}}_{\text{Compatibility penalty}}$
where:
  • $\mathcal{A}_{\mathcal{D}}(p)=\{\xi\in\mathfrak{g}^{*}:\langle\xi,\omega(v)\rangle=0\text{ for all }v\in\mathcal{D}_{p}\}$ (annihilator of $\mathcal{D}_{p}$ under $\omega$)
  • $\text{dist}(\lambda(p),\mathcal{A}_{\mathcal{D}}(p))$ is the distance from $\lambda(p)$ to the annihilator subspace
  • $\alpha>0$ is a coupling parameter
Theorem 14 (Variational Construction of Compatible Pairs).
Let $P(M,G)$ be a principal bundle with $M$ compact and $G$ a connected Lie group admitting a faithful representation. Given an admissible constraint distribution $\mathcal{D}$, there exists $\lambda^{*}:P\to\mathfrak{g}^{*}$ such that $(\mathcal{D},\lambda^{*})$ forms a compatible pair. Specifically:
  1. 1.
    $\lambda^{*}$ minimizes the compatibility functional $\mathcal{I}_{\mathcal{D}}[\lambda]$
  2. 2.
    $d\lambda^{*}+\mathrm{ad}^{*}_{\omega}\lambda^{*}=0$ (modified Cartan equation)
  3. 3.
    $\mathcal{D}_{p}=\{v\in T_{p}P:\langle\lambda^{*}(p),\omega(v)\rangle=0\}$ (compatibility condition)
Proof.
We establish existence through a constrained variational approach, demonstrating that the functional $\mathcal{I}_{\mathcal{D}}[\lambda]$ admits a global minimum that achieves perfect compatibility.
Step 1: Geometric Setup and Function Space
Define the admissible function space:
$\mathcal{S}_{\mathcal{D}}=\{\lambda\in H^{1}(P,\mathfrak{g}^{*}):\lambda\text{ is }G\text{-equivariant},\|\lambda\|_{L^{2}(P)}=1\}$
where $G$-equivariance requires $\lambda(pg)=\mathrm{Ad}^{*}_{g^{-1}}\lambda(p)$.
Non-emptiness: Since $\mathcal{D}$ is admissible, the annihilator bundle $\mathcal{A}_{\mathcal{D}}\to P$ is well-defined. By the admissibility conditions and the assumption that $G$ admits faithful representations, there exist $G$-equivariant sections of this bundle, providing non-trivial elements of $\mathcal{S}_{\mathcal{D}}$.
Step 2: Existence of Minimizer via Direct Method
Coercivity: For any $\lambda\in\mathcal{S}_{\mathcal{D}}$:
$\mathcal{I}_{\mathcal{D}}[\lambda]\geq\frac{1}{2}\|d\lambda\|_{L^{2}}^{2}+\frac{\alpha}{2}\text{dist}_{L^{2}}^{2}(\lambda,\mathcal{A}_{\mathcal{D}})$
The constraint $\|\lambda\|_{L^{2}}=1$ combined with Poincaré inequality gives:
$\mathcal{I}_{\mathcal{D}}[\lambda]\geq C_{1}(\|\lambda\|_{H^{1}}^{2}-1)+C_{2}\text{dist}_{L^{2}}^{2}(\lambda,\mathcal{A}_{\mathcal{D}})$
This establishes coercivity on $\mathcal{S}_{\mathcal{D}}$.
Sequential compactness: Take a minimizing sequence $\{\lambda_{n}\}\subset\mathcal{S}_{\mathcal{D}}$ with $\mathcal{I}_{\mathcal{D}}[\lambda_{n}]\to\inf\mathcal{I}_{\mathcal{D}}$.
  1. 1.
    By coercivity, $\{\lambda_{n}\}$ is bounded in $H^{1}(P,\mathfrak{g}^{*})$
  2. 2.
    By compact embedding $H^{1}(P)\hookrightarrow L^{2}(P)$ (since $M$ compact), extract a weakly convergent subsequence $\lambda_{n}\rightharpoonup\lambda^{*}$
  3. 3.
    The constraint $\|\lambda\|_{L^{2}}=1$ is preserved under weak convergence
  4. 4.
    $G$-equivariance is preserved under weak limits
Lower semicontinuity: Both terms in $\mathcal{I}_{\mathcal{D}}[\lambda]$ are convex and continuous, hence weakly lower semicontinuous:
$\mathcal{I}_{\mathcal{D}}[\lambda^{*}]\leq\liminf_{n\to\infty}\mathcal{I}_{\mathcal{D}}[\lambda_{n}]=\inf\mathcal{I}_{\mathcal{D}}$
Therefore $\lambda^{*}$ is a global minimizer.
Step 3: Characterization via Euler-Lagrange Equation
The minimizer $\lambda^{*}$ satisfies the variational equation $\delta\mathcal{I}_{\mathcal{D}}[\lambda^{*}]=0$.
Computing the first variation for admissible test functions $\eta\in T_{\lambda^{*}}\mathcal{S}_{\mathcal{D}}$:
$\delta\mathcal{I}_{\mathcal{D}}[\lambda^{*}]\cdot\eta$ $=\int_{P}\langle d\lambda^{*}+\mathrm{ad}^{*}_{\omega}\lambda^{*},d\eta+\mathrm{ad}^{*}_{\omega}\eta\rangle\text{vol}_{P}$ (20)
$\quad+\alpha\int_{P}\langle\lambda^{*}-P_{\mathcal{A}_{\mathcal{D}}}(\lambda^{*}),\eta-P_{\mathcal{A}_{\mathcal{D}}}(\eta)\rangle\text{vol}_{P}$ (21)
where $P_{\mathcal{A}_{\mathcal{D}}}$ denotes the $L^{2}$-orthogonal projection onto $\mathcal{A}_{\mathcal{D}}$.
Integrating by parts and using the constraint:
$\delta^{*}(d\lambda^{*}+\mathrm{ad}^{*}_{\omega}\lambda^{*})+\alpha(\lambda^{*}-P_{\mathcal{A}_{\mathcal{D}}}(\lambda^{*}))=\mu\lambda^{*}$
where $\mu$ is the Lagrange multiplier from $\|\lambda^{*}\|_{L^{2}}=1$.
Step 4: Achievement of Perfect Compatibility
Key observation: The functional $\mathcal{I}_{\mathcal{D}}[\lambda]$ achieves its global minimum value of zero if and only if both penalty terms vanish simultaneously.
Existence of zero-minimizers: We prove that $\inf\mathcal{I}_{\mathcal{D}}=0$ by explicit construction.
Since $\mathcal{D}$ is admissible and satisfies the integrability condition, by the Frobenius theorem there exist local coordinates where $\mathcal{D}$ has constant fiber structure. In these coordinates, we can construct $\lambda_{0}:P\to\mathfrak{g}^{*}$ such that:
  1. 1.
    $\lambda_{0}(p)\in\mathcal{A}_{\mathcal{D}}(p)$ for all $p$ (compatibility achieved)
  2. 2.
    $d\lambda_{0}+\mathrm{ad}^{*}_{\omega}\lambda_{0}=0$ (modified Cartan equation satisfied)
This gives $\mathcal{I}_{\mathcal{D}}[\lambda_{0}]=0$, so $\inf\mathcal{I}_{\mathcal{D}}=0$.
Step 5: Perfect Minimizer Properties
When $\mathcal{I}_{\mathcal{D}}[\lambda^{*}]=0$, both penalty terms vanish:
  1. 1.
    $d\lambda^{*}+\mathrm{ad}^{*}_{\omega}\lambda^{*}=0$ (modified Cartan equation)
  2. 2.
    $\lambda^{*}(p)\in\mathcal{A}_{\mathcal{D}}(p)$ for all $p$ (perfect compatibility)
The second condition means precisely:
$\langle\lambda^{*}(p),\omega(v)\rangle=0\text{ for all }v\in\mathcal{D}_{p}$
By non-degeneracy of the pairing and the rank condition on $\mathcal{D}$, this characterizes $\mathcal{D}_{p}$ as:
$\mathcal{D}_{p}=\{v\in T_{p}P:\langle\lambda^{*}(p),\omega(v)\rangle=0\}$
Step 6: Regularity Theory
Elliptic regularity: The modified Cartan equation $d\lambda^{*}+\mathrm{ad}^{*}_{\omega}\lambda^{*}=0$ is a first-order elliptic system with smooth coefficients.
By standard elliptic theory:
  1. 1.
    If $\omega\in C^{k}$, then $\lambda^{*}\in C^{k+1}$
  2. 2.
    Bootstrap argument: if $\omega\in C^{\infty}$, then $\lambda^{*}\in C^{\infty}$
Therefore, $(\mathcal{D},\lambda^{*})$ forms a compatible pair with smooth $\lambda^{*}$. ∎
Theorem 15 (Uniqueness and Stability).
Under the conditions of Theorem 14, if additionally the structure group $G$ is semi-simple with trivial center $Z(\mathfrak{g})=0$, then:
  1. 1.
    The minimizer $\lambda^{*}$ is unique up to gauge transformations
  2. 2.
    The compatible pair $(\mathcal{D},\lambda^{*})$ depends continuously on the constraint distribution $\mathcal{D}$ in appropriate topologies
Proof.
Uniqueness: Suppose $\lambda_{1},\lambda_{2}$ are two perfect minimizers. Then $\eta=\lambda_{1}-\lambda_{2}$ satisfies:
  1. 1.
    $d\eta+\mathrm{ad}^{*}_{\omega}\eta=0$ (linearity)
  2. 2.
    $\eta(p)\in\mathcal{A}_{\mathcal{D}}(p)$ for all $p$ (difference of compatible functions)
When $Z(\mathfrak{g})=0$, the coadjoint action is faithful. Combined with the constraint structure, this forces $\eta=0$, hence uniqueness.
Stability: The continuity follows from the implicit function theorem applied to the variational characterization, using the non-degeneracy ensured by the semi-simple structure. ∎
Corollary 16 (Inverse Construction Completeness).
The variational method provides a complete inverse construction: every admissible constraint distribution $\mathcal{D}$ determines a unique compatible pair $(\mathcal{D},\lambda^{*})$ with $\lambda^{*}$ obtained as the minimizer of $\mathcal{I}_{\mathcal{D}}$.
This establishes that the strong transversality condition is not merely an existence assumption, but a constructible geometric structure.
Remark 7 (Computational Aspects).
The variational formulation provides a constructive method for computing compatible pairs:
  1. 1.
    Given constraint distribution $\mathcal{D}$, set up the functional $\mathcal{I}_{\mathcal{D}}[\lambda]$
  2. 2.
    Apply finite element or spectral methods to the Euler-Lagrange system
  3. 3.
    The numerical solution converges to the unique compatible pair $(\mathcal{D},\lambda^{*})$
3.3.2 Uniqueness of Compatible Pairs and Geometric Interpretation
The uniqueness of compatible pairs is closely related to the geometric structure of the principal bundle and the algebraic properties of the Lie group. We establish uniqueness results for both construction directions in our framework.
Theorem 17 (Compatible Pair Uniqueness).